# LLMGA: Multimodal Large Language Model based Generation Assistant

### 10 Dec 2023 v2

@@ Abstract
In this paper, we introduce a Multimodal Large Language Model-based Generation Assistant (LLMGA), leveraging the vast reservoir of knowledge and proficiency in reasoning, comprehension, and response inherent in Large Language Models (LLMs) to assist users in image generation and editing. Diverging from existing approaches where Multimodal Large Language Models (MLLMs) generate fixed-size embeddings to control Stable Diffusion (SD), our LLMGA provides a detailed language generation prompt for precise control over SD. This not only augments LLM context understanding but also reduces noise in generation prompts, yields images with more intricate and precise content, and elevates the interpretability of the network. To this end, we curate a comprehensive dataset comprising prompt refinement, similar image generation, inpainting &\&& outpainting, and visual question answering. Moreover, we propose a two-stage training scheme. In the first stage, we train the MLLM to grasp the properties of image generation and editing, enabling it to generate detailed prompts. In the second stage, we optimize SD to align with the MLLM’s generation prompts. Additionally, we propose a reference-based restoration network to alleviate texture, brightness, and contrast disparities between generated and preserved regions during image editing. Extensive results show that LLMGA has promising generative capabilities and can enable wider applications in an interactive manner.
@@ 1 Introduction
Recent studies have begun delving deeper into Multimodal Large Language Models (MLLMs) [3, 67] built upon LLMs, aiming to empower LLMs to comprehend inputs extending beyond text. For example, BLIP-2 [29] and LLaVA [31] employ visual encoders to transform images into input embeddings, enabling them to be used as prompts alongside text input within the LLM, thus achieving compatibility of LLMs with the visual modality. Furthermore, recent works focused on extending the capabilities of LLMs to generate multimodal outputs. For example, GILL [28] involves instructing LLMs to predict fixed-size visual embeddings aligned with CLIP [38] space to control the Stable Diffusion [40] (SD) for image generation.
Figure 1: Examples of LLMGA for assisting in image generation and editing. (1) For T2I generation, LLMGA can refine the user’s generation prompt to produce more vivid and vibrant images. (2) For similar image generation, LLMGA can understand the component and layout of the input images and generate a similar image. (3) For inpainting and outpainting, LLMGA can provide detailed generation prompts based on user preferences and input images. (4) Interactive image generation and editing exemplify the comprehensive capabilities of LLMGA. Users can design satisfactory images by engaging in interactions with LLMGA, leveraging its vast knowledge and ideas.
@@
[…]
Moreover, we have identified noticeable disparities in texture, contrast, and brightness between the newly generated and preserved sections in SD-based image editing, such as inpainting. To address these issues, we propose a Diffusion-based Reference Restoration Network (DiffRIR), based on the structure of the SOTA restoration network DiffIR [59]. Specifically, aside from images generated by SD, we add masked images as reference inputs into DiffRIR. This enables the DiffRIR to refer to the texture, contrast, and brightness of the preserved regions for restoration. Additionally, we introduce perturbations to contrast and brightness during training, making DiffRIR correct contrast and brightness disparities in the images.
As shown Fig. 1, the remarkable capabilities of the LLMGA come to light, leaving us in awe. LLMGA excels in a myriad of functions, elevating image generation and editing, realizing interactive image editing, and delivering astonishing capabilities for image design due to its expansive world knowledge and robust reasoning abilities.
In summary, our contributions are as follows:
  • We introduced LLMGA, a simple, powerful, and generalizable multimodal generation model. Extensive experiments affirm the efficacy of LLMGA in enhancing image generation and editing through its vast knowledge base.
  • We construct a training dataset, including four parts: prompt refinement, similar image generation, inpainting & outpainting, and visual question answering. This enhances LLMGA’s comprehension of generation and editing tasks while standardizing response formats.
@@
  • Open-source. The following assets are released: the generated data, the codebase for data generation and model training, the model checkpoint, and a demo.
@@ 2 Related Work
Diffusion Model. Diffusion Models (DMs) [48, 22, 49, 27, 14, 23, 4, 20, 33, 36, 6] have achieved remarkable results in image generation. DMs adopt a parameterized Markov chain to optimize the lower variational bound on the likelihood function. In this way, it can generate realistic images from Gaussian noise. After that, several DM methods [59, 58, 32, 42, 7, 17, 25, 26, 54, 8, 37, 61, 18, 5, 44, 13] have been tailored to enhance the text-to-image (T2I) generation and editing. Notably, GLIDE [34] pioneered the incorporation of text features into transformer blocks during the denoising process. Subsequently, DALL-E [39], Imagen [43], and Stable Diffusion [40] have made substantial strides in improving T2I generation. Subsequently, some works  [64, 41, 65] introduced conditioning controls to the DMs to facilitate a more convenient and precise manipulation of the generation process. Overall, enhancing the user-friendliness of DMs is a key focus within the community. In this paper, we introduce LLMGA, leveraging the extensive knowledge and powerful reasoning capabilities of LLM to facilitate users in achieving more easily attainable and satisfactory image designs.
@@
Afterward, the community began focusing on the development of the Multimodal Large Language Model [67, 24, 16, 62, 15, 19, 63, 2]. They aim to enable LLMs to comprehend both images and text and provide textual responses. For instance, Flamingo [3] encodes images and feeds them into the LLM’s attention layer. BLIP-2 [29] employs Q-Former to encode input images into queries. Furthermore, LLaVA [31] utilizes CLIP [38] to encode images into image embeddings and concatenate them with text embeddings.
Recent concurrent works, such as Next-GPT [57], have extended the capabilities to encompass audio and video modalities. Moreover, Visual-ChatGPT [56] and HuggingGPT [47] make LLMs act as agents capable of invoking various pre-trained visual models to achieve MLLM. However, these works focus on making LLM determine the combined invocation of modules (such as detection, recognition, and generation) to fulfill user requirements. However, these methods are not tailored for generation and editing and use concise prompts that lack the capability to enhance results. Thus, we propose LLMGA, which is designed to assist with image generation and editing. It can achieve satisfactory results by reasoning and interaction with the user.
@@ 3 Methodology
Figure 2: The overview of LLMGA. (a) In the first training stage, we train the MLLM to produce generation prompts based on provided instructions. Moreover, we construct a training dataset including four categories: prompt refinement, similar image generation, inpainting & outpainting, and visual question answering. (b) In the second training stage, we optimize SD to be adapt to the detailed generation prompts from MLLM. (c) In the inference stage, LLMGA can respond to user queries and assist in image generation and editing.
@@ 3.1 Overview of LLMGA
However, we need to address several issues: (1) As a language assistant, MLLM may decline the execution of generation instructions. (2) MLLM lacks a nuanced understanding of image generation and editing, and cannot produce precise and detailed generation prompts. (3) MLLM needs to decide which part of the output text serves as generation prompts to guide generation. (4) SD’s CLIP encode only $75$ tokens. Moreover, SD is trained on short captions, while detailed prompts generated by LLMGA may exceed $150$ tokens.This disparity makes it hard for SD to understand the instructions from LLMGA. To this end, we construct a training dataset and two-stage training schemes, which train the MLLM (Sec. 3.2) and SD (Sec. 3.3).
@@
where $\mathbf{T}_{input}$ indicates the input text instructions. Notably, $\mathcal{F}_{MLLM}$ can process only $\mathbf{T}_{input}$ as input.
[…]
Next, we output $\mathbf{T}_{R}$ directly, while $\mathbf{T}_{P}$ is fed into the subsequent SD for T2I generation (Eq. 2) or editing (Eq. 3). Notably, for image editing such as inpainting and outpainting, an additional mask $\mathbf{I}_{mask}$ and the edited image $\mathbf{I}_{input}$ are essential inputs for the inpainting SD. These inputs specify the region requiring editing.
$\vspace{-1mm}\mathbf{I}_{T2I}=\mathcal{F}_{T2I}(\mathbf{T}_{P},\mathbf{Z}),$ (2)
$\vspace{-1mm}\mathbf{I}_{edit}=\mathcal{F}_{edit}(\mathbf{I}_{input},\mathbf{I}_{mask},\mathbf{T}_{P},\mathbf{Z}),$ (3)
Where $\mathbf{Z}$ denotes the random Gaussian noise. Furthermore, to ensure the encoding of all $\mathbf{T}_{P}$ for SD, we iteratively run the CLIP text encoder until all prompts are encoded.
@@ 3.2 MLLM Training
As described in Sec. 3.1, original MLLMs are specifically designed and trained as language assistants, but they lack the proficiency to assist in image generation and editing. Notably, considering the input of images and the performance of open-sourced LLMs, employing few-shot learning to guide the model to achieve the desired results proves challenging and inefficient. Therefore, it is crucial to train the MLLM to be a generation assistant, know the response format, and elevate its comprehension of the properties of image generation and editing. Thus, as shown in Fig. 2(a), we construct a training dataset consisting of four categories:
[…]
    Figure 3: Visual comparison on T2I. LLMGA can refine short prompts by adding details, such as clothing, background, and actions.
  • Visual Question Answering. To preserve the Visual Question Answering (VQA) capability of MLLM, we have integrated the VQA dataset into the training of LLMGA. Specifically, we utilized the LLaVA-Instruct-150k dataset [31], which was generated by GPT-4.
@@ 3.3 Stable Diffusion Training
As described in Sec. 3.1, the original SD’s CLIP text encoder only encodes $75$ tokens, which cannot handle the entire MLLM’s generation prompt. Moreover, the original SD is trained on brief captions, which cannot fully understand the generation prompts. Thus, we repeatedly use the VAE text encoder to encode all instances of $\mathbf{T}_{P}$ for SD. Besides, we train the T2I SD model and the inpainting SD model, respectively. For both generation and editing tasks, the generation prompts of MLLM are detailed descriptions of images. Therefore, during training, we can instruct MLLM to provide a detailed description $\mathbf{T}_{P}$ for images from the LAION-Aesthetics and MSCOCO datasets. Subsequently, $\mathbf{T}_{P}$ is fed into T2I SD or inpainting SD for joint training. Notably, we only optimize the SD unet while freezing the parameters of other networks. To accelerate the training process, we record the prompt $\mathbf{T}_{P}$ of MLLM to avoid redundant calculations. The model is trained using SD loss:
@@
Table 1: Quantitative comparison on T2I generation on the subset of MSCOCO dataset. The best and second best performances are marked in bold and underlined, respectively.
Method FIDnormal-$\downarrow$ ISnormal-$\uparrow$ CLIP-Inormal-$\uparrow$
SD1.5 142.86 13.50 68.93
SD1.5-ft 141.99 13.57 71.05
GILL [28] 150.07 12.19 67.63
LLMGA-7B (SD1.5) 147.13 13.39 70.21
LLMGA-7B (SD1.5-ft, Ours) 139.46 13.63 72.26
LLMGA-13b (SD1.5) 145.97 13.41 70.24
LLMGA-13b (SD1.5-ft, Ours) 136.23 13.69 73.24
Figure 4: Visual comparison on inpainting and outpainting. LLMGA can infer complete images based on input masked images.
@@ 3.4 Restoration Network Training
For SD inpainting and outpainting, we observed noticeable disparities between the preserved and newly generated regions in the edited images. These disparities can be attributed to variations in texture, contrast, and brightness. Besides, after finishing image generation and editing, it is customary to employ image restoration methods [59, 58, 55, 60] on the generated images to enhance the quality. Therefore, we intend to employ a restoration network to address these disparities in the edited images.
To enhance the consistency between the newly generated and the preserved regions, we introduce a reference-based restoration scheme. Specifically, existing restoration methods [59, 55, 60] take low-quality (LQ) images as input and produce high-quality (HQ) images, but they often do not leverage the preserved information from the given masked image. Different from these restoration methods, we concatenate the LQ image ${I}_{LQ}$ and the masked image, i.e., $(1-\mathbf{I}_{mask})\mathbf{I}_{GT}$, as inputs and feed it into the restoration model $\mathcal{F}_{R}$. This process can be formulated as:
@@
$\vspace{-2mm}\mathcal{D}_{2}(\mathbf{x})=c_{1}\mathbf{x}+c_{2},$ (7)
@@
where $\mathcal{D}_{2}$ indicates contrast and brightness disturbance, $\mathbf{x}$ denotes the input image. $c_{1}$ is the contrast gain, randomly varied within the range of $[0.94,1.06]$, while $c_{2}$ is the brightness bias, randomly varied within the range of $[-0.05,0.05]$. $\mathcal{D}_{1}$ represents the real-world degradation process used in restoration model [55, 59, 60]. $\mathbf{I}_{GT}$ is ground truth image, and $\mathbf{I}_{LQ}$ is LQ image. In this paper, we adopt the network structure of the SOTA restoration model DiffIR [59] and apply our schemes to it to obtain our Diffusion-based Reference Restoration Network (DiffRIR).

### 11 Mar 2024 v3

# https://llmga.github.io/

@@ Abstract
In this paper, we introduce a Multimodal Large Language Model-based Generation Assistant (LLMGA), leveraging the vast reservoir of knowledge and proficiency in reasoning, comprehension, and response inherent in Large Language Models (LLMs) to assist users in image generation and editing. Diverging from existing approaches where Multimodal Large Language Models (MLLMs) generate fixed-size embeddings to control Stable Diffusion (SD), our LLMGA provides a detailed language generation prompt for precise control over SD. This not only augments LLM context understanding but also reduces noise in generation prompts, yields images with more intricate and precise content, and elevates the interpretability of the network. To this end, we curate a comprehensive dataset comprising prompt refinement, similar image generation, inpainting & outpainting, and instruction-based editing. Moreover, we propose a two-stage training scheme. In the first stage, we train the MLLM to grasp the properties of image generation and editing, enabling it to generate detailed prompts. In the second stage, we optimize SD to align with the MLLM’s generation prompts. Additionally, we propose a reference-based restoration network to alleviate texture, brightness, and contrast disparities between generated and preserved regions during inpainting and outpainting. Extensive results show that LLMGA has promising generation and editing capabilities and can enable more flexible and expansive applications in an interactive manner.
Keywords:
Interactive generation and editing LLM Diffusion Model
@@ 1 Introduction
Recent studies have begun delving deeper into Multimodal Large Language Models (MLLMs) [2, 71] built upon LLMs, aiming to empower LLMs to comprehend inputs extending beyond text. For example, BLIP-2 [31] and LLaVA [34] employ visual encoders to transform images into input embeddings, enabling them to be used as prompts alongside text input for the LLM, thus achieving LLMs with the visual modality. Furthermore, recent works focused on extending the capabilities of LLMs to generate multimodal outputs. For example, GILL [30] involves instructing LLMs to predict fixed-size visual embeddings aligned with CLIP [41] space to control the Stable Diffusion [43] (SD) for image generation.
Figure 1: Some examples of LLMGA for assisting in image generation and editing. (1) T2I generation. LLMGA can refine the user’s generation prompt to produce more vivid and vibrant images. (2) Similar image generation. LLMGA can understand the component and layout of the input images and generate a similar image. (3) Inpainting & Outpainting. LLMGA can provide detailed generation prompts based on user preferences and input images. (4) Instruction based editing. LLMGA can understand user instructions and realize accurate editing. (5) Interactive image generation and editing exemplify the comprehensive capabilities of LLMGA. Users can design satisfactory images by engaging in interactions with LLMGA, leveraging its vast knowledge.
@@
To this end, we have devised a two-stage training scheme. First, we construct a training dataset: prompt refinement, similar image generation, inpainting & outpainting, and visual question answering. We then train LLMGA on these four datasets to cultivate four fundamental capabilities: (1) For concise user prompts, LLMGA can refine the generation of intricate details, encompassing attire, background, and characters. (2) LLMGA can precisely regenerate an image it observes. (3) LLMGA can generate or refine prompts for inpainting & outpainting based on its understanding of the image. (4) LLMGA can generate accurate prompts for instruction-based editing according to users’ requirements and given images. Additionally, we make LLMGA use special symbols <gen_img> and </gen_img> to distinguish generation prompts and responses. In the second stage, we freeze the parameters of LLMGA’s MLLM and initiate joint training with the SD. This process enables the SD to acclimate to the detailed prompt produced by the MLLM. Notably, when the input token count exceeds $75$, we iteratively apply the CLIP [41] encoder to the surplus tokens.
Moreover, we have identified noticeable disparities in texture, contrast, and brightness between the newly generated and preserved sections in inpainting & outpainting. Therefore, we propose a Diffusion-based Reference Restoration Network (DiffRIR). Specifically, aside from images generated by SD, we add masked images as reference inputs into DiffRIR. This enables the DiffRIR to refer to the texture, contrast, and brightness of the preserved regions for restoration. Additionally, we introduce perturbations to contrast and brightness during training, enabling DiffRIR to correct contrast and brightness disparities in the images.
As shown in Fig. 1, LLMGA is a unified and interactive framework for image generation and editing, endowed with a wide array of capabilities: (1) LLMGA leverages its extensive world knowledge and powerful reasoning abilities to assist image generation and editing and significantly improve results. (2) LLMGA can be integrated with external plugins, like ControlNet. (3) Most importantly, users can interact with LLMGA to design satisfying images in a more convenient, flexible, and enjoyable way. In summary, our contributions are as follows:
  • We proposed LLMGA, a simple yet powerful interactive generation and editing framework. Experiments affirm the efficacy of LLMGA in enhancing generation and editing thanks to its vast knowledge and interactive features. Plus, LLMGA can integrate with external plugins for wider applications.
  • We construct a training dataset, including four parts: prompt refinement, similar image generation, inpainting & outpainting, and instruction-based editing. This enhances LLMGA’s comprehension of generation and editing tasks while standardizing response formats.
@@
  • Open-source. The following assets are released: the generated data, the codebase for model training, the model checkpoint, and a demo.
@@ 2 Related Work
Diffusion Model. Diffusion Models (DMs) [50, 22, 52, 29, 14, 23, 3, 20, 36, 39, 5, 8] have achieved remarkable results in image generation. DMs adopt a parameterized Markov chain to optimize the lower variational bound on the likelihood function. In this way, it can generate realistic images from Gaussian noise. After that, several DM methods [62, 61, 35, 45, 6, 17, 27, 28, 57, 7, 40, 64, 18, 4, 46, 13] have been tailored to enhance the text-to-image (T2I) generation and editing. Notably, GLIDE [37] pioneered the incorporation of text features into transformer blocks during the denoising process. Subsequently, DALL-E [42], Imagen [47], and Stable Diffusion [43] have made substantial strides in improving T2I generation. Subsequently, some works  [68, 44, 69] introduced conditioning controls to the DMs to facilitate a more convenient and precise manipulation of the generation process. Overall, enhancing the user-friendliness of DMs is a key focus within the community. In this paper, we introduce LLMGA, leveraging the extensive knowledge and powerful reasoning capabilities of LLM to facilitate users in achieving more easily attainable and satisfactory image designs.
@@
Afterward, the community began focusing on the development of the Multimodal Large Language Model [71, 24, 16, 65, 15, 19, 66]. They aim to enable LLMs to comprehend both images and text and provide textual responses. For instance, Flamingo [2] encodes images and feeds them into the LLM’s attention layer. BLIP-2 [31] employs Q-Former to encode input images into queries. Additionally, LLaVA [34] leverages CLIP [41] to encode images into visual embeddings, which are then concatenated with text embeddings.
Recent concurrent works, such as Next-GPT [60], have extended the capabilities to encompass audio and video modalities. Moreover, Visual-ChatGPT [59] and HuggingGPT [49] make LLMs act as agents capable of invoking various pre-trained visual models to achieve MLLM. However, these works focus on making LLM determine the combined invocation of modules (such as detection, recognition, and generation) to fulfill user requirements. However, these methods are not tailored for generation and editing and use concise prompts that lack the capability to enhance results. Thus, we propose LLMGA, which is designed to assist with various image generation and editing tasks. It can achieve satisfactory results by strong reasoning capability and flexible interaction with users.
@@ 3 Methodology
Figure 2: The overview of LLMGA. (a) In the first training stage, we train the MLLM to produce generation prompts based on provided instructions. Moreover, we construct a training dataset including four categories: prompt refinement, similar image generation, inpainting & outpainting, and instruction-based editing. (b) In the second training stage, we optimize SD to adapt to the detailed generation prompts from MLLM. (c) In the inference stage, LLMGA can respond to user queries and assist in various tasks, such as image generation, inpainting & outpainting, and editing.
@@ 3.1 Overview of LLMGA
However, we need to address several issues: (1) As a language assistant, MLLM may decline the execution of generation instructions. (2) MLLM lacks a nuanced understanding of image generation and editing, and cannot produce precise and detailed generation prompts. (3) MLLM needs to decide which part of the output text serves as generation prompts to guide generation. (4) SD’s CLIP encode only $75$ tokens. Moreover, SD is primarily trained on short captions, while detailed prompts generated by LLMGA may exceed $150$ tokens. This disparity makes it hard for SD to understand the instructions from LLMGA. To address the aforementioned challenges, we construct a training dataset and design two-stage training schemes to train the MLLM (Sec. 3.2) and SD (Sec. 3.3).
@@
where $\mathbf{T}_{input}$ indicates the input text instructions from users. It is notable that $\mathcal{F}_{MLLM}$ can process only $\mathbf{T}_{input}$ as input.
The text output $\mathbf{T}_{output}$ can comprise two components: text response $\mathbf{T}_{R}$ and generation prompt $\mathbf{T}_{P}$. To distinguish between $\mathbf{T}_{R}$ and $\mathbf{T}_{P}$, we adopt new special tokens, i.e., <gen_img> and </gen_img>, to encompass $\mathbf{T}_{P}$.
We present $\mathbf{T}{R}$ as the immediate text response to users. Concurrently, $\mathbf{T}_{P}$ is further fed into the subsequent SD to guide T2I generation (Eq. 2), inpainting & outpainting (Eq. 3), and instruction-based image editing (Eq. 4).
$\vspace{-1mm}\mathbf{I}_{g}=\mathcal{F}_{T2I}(\mathbf{T}_{P},\mathbf{Z}),$ (2)
$\vspace{-1mm}\mathbf{I}_{p}=\mathcal{F}_{Paint}(\mathbf{I}_{input},\mathbf{I}_{mask},\mathbf{T}_{P},\mathbf{Z}),$ (3)
$\mathbf{I}_{e}=\mathcal{F}_{Edit}(\mathbf{I}_{input},\mathbf{T}_{P}),$ (4)
Where $\mathbf{Z}$ denotes the random Gaussian noise. Furthermore, to ensure the encoding of all $\mathbf{T}_{P}$ for SD, we iteratively run the CLIP text encoder until all prompts are encoded. For inpainting & outpainting, except the input image $\mathbf{I}_{input}$, an additional mask $\mathbf{I}_{mask}$ are essential inputs for the inpainting SD to specify the region requiring generation. For instruction-based image editing, $\mathcal{F}_{Edit}$ performs inversion [51, 26] and prompt-to-prompt [21] based on T2I SD.
@@ 3.2 MLLM Training
As described in Sec. 3.1, original MLLMs are specifically designed and trained as language assistants, but they lack the proficiency to assist in image generation, and editing. Notably, considering the input of images and the performance of open-sourced MLLMs, employing few-shot learning to guide the model to achieve the desired results proves challenging and inefficient. Therefore, it is crucial to train the MLLM to serve as proficient assistants in image generation and editing tasks, understanding the expected response formats and enhancing their comprehension of image generation and editing properties. To this end, as depicted in Fig. 2 (a), we construct a training dataset consisting of four categories:
  • Prompt Refinement. We establish this dataset to cultivate the prompt refinement ability of the MLLM. Specifically, we utilize GPT4-V to furnish detailed descriptions of images in MSCOCO [32]. These detailed descriptions, along with the original MSCOCO brief descriptions, constitute a training text pair. During training, we input the brief MSCOCO captions and randomly select and append a generation instruction, such as “Generate this image”, or a description instruction, like "Describe this sentence in detail”. When a generation instruction is included in the prompt, we add <gen_img> and </gen_img> on the later detailed description for training.
  • Similar Image Generation. We select images from the MSCOCO dataset along with corresponding detailed descriptions generated by GPT4-V to create the Similar Image Generation dataset. During training, we input the images along with a generation instruction, such as “Generate a similar image.” or a description instruction, such as “Describe the following image.”. In cases where a generation instruction is provided, we add <gen_img> and </gen_img> on the subsequent detailed description for training.
  • Inpainting &\&& Outpainting. We use pairs of detailed descriptions and images from the Similar Image Generation dataset. During training, we input masked images with inpainting or outpainting instructions, like “Inpaint this image.” or “Outpaint this image.”. Besides, we include <gen_img> and </gen_img> on the subsequent detailed description during training.
  • Instruction-based editing. We fine-tune Mixtral-8x7B [25] to enable it to automatically generate editing data based on detailed descriptions from MSCOCO. For example, given the original caption “A cat is lying on the ground”, Mixtral-8x7B generates the output “Instruction: Replace the cat with a dog., Target caption: A dog is lying on the ground”. Subsequently, we clean the generated data. During training, we can input MSCOCO images or original captions and provide corresponding editing instructions, aiming to train LLMGA to output the target caption. Additionally, we include <gen_img> and </gen_img> tags on the target caption during training.
During training, alongside the data we curated, we integrated the Alpaca [54] dataset regenerated by GPT4, to enhance LLMGA’s question-answering (QA). Notably, we excluded certain visual multimodal incompatible question-answer pairs from it. Furthermore, we incorporated the LLaVA v1.5 mix665k dataset [33] to endow LLMGA with Visual Question Answering (VQA) capabilities.
@@ 3.3 Stable Diffusion Training
As described in Sec. 3.1, the original SD’s CLIP text encoder only encodes $75$ tokens, which cannot handle the entire MLLM’s generation prompt. Moreover, the original SD is trained on brief captions, which cannot fully understand the generation prompts. Thus, we repeatedly use the CLIP text encoder to encode all instances of $\mathbf{T}_{P}$ for SD. Besides, we train the T2I SD model and the inpainting SD model, respectively. For both generation and inpainting & outpainting tasks, the generation prompts of MLLM are detailed descriptions of images. Therefore, during training, we can instruct MLLM to provide a detailed description $\mathbf{T}_{P}$ for images from the LAION-Aesthetics [48] and MSCOCO datasets. Subsequently, $\mathbf{T}_{P}$ is fed into T2I SD or inpainting SD for joint training. Notably, we only optimize the SD unet while freezing the parameters of other networks. To accelerate the training, we record the prompt $\mathbf{T}_{P}$ of MLLM to avoid redundant calculations. The model is trained using SD loss (Eq. 6). For instruction-based editing, we directly adopt the pre-trained T2I SD model.
@@ 3.4 Restoration Network Training
For SD inpainting & outpainting, we observed noticeable disparities between the preserved and newly generated regions in the edited images. These disparities can be attributed to variations in texture, contrast, and brightness. Besides, after finishing image generation and editing, it is customary to employ image restoration methods [62, 61, 58, 63] on the generated images to enhance the quality. Therefore, we intend to employ an image restoration network to address these disparities in the inpainting & outpainting images.
To enhance the consistency between the newly generated and the preserved regions, we introduce a reference-based restoration scheme. Specifically, existing restoration methods [62, 58, 63] take low-quality (LQ) images as input and produce high-quality (HQ) images, but they often do not leverage the preserved information from the given masked image. Different from them, we concatenate the LQ image ${I}_{LQ}$ and the masked image, i.e., $(1-\mathbf{I}_{mask})\mathbf{I}_{GT}$, as inputs and feed them into the restoration model $\mathcal{F}_{R}$. This process can be formulated as:
@@
$\vspace{-5mm}\mathcal{D}_{2}(\mathbf{x})=c_{1}\mathbf{x}+c_{2},$ (8)
@@
where $\mathcal{D}_{2}$ indicates contrast and brightness disturbance, $\mathbf{x}$ denotes the input image. $c_{1}$ is the contrast gain, randomly varied within the range of $[0.94,1.06]$, while $c_{2}$ is the brightness bias, randomly varied within the range of $[-0.05,0.05]$. $\mathcal{D}_{1}$ represents the real-world degradation process used in restoration model [58, 62, 63]. $\mathbf{I}_{GT}$ is ground truth image, and $\mathbf{I}_{LQ}$ is LQ image. Here, we adopt the network structure of the SOTA restoration model DiffIR [62] and apply our schemes to it to obtain our Diffusion-based Reference Restoration Network (DiffRIR).